OpenStudy (anonymous):

@rvc @kelliegirl33 how do i factor 3x^2-4x-7

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

This video helped me @knightmare6 http://www.virtualnerd.com/algebra-1/polynomials-and-factoring/trinomial-factorization-example.php

OpenStudy (michele_laino):

hint: the requested factorization, is given by the subsequent formula: \[\Large 3\left( {x - {x_1}} \right)\left( {x - {x_2}} \right)\] where x_1 and x_2 are the solution of this equation: \[\Large 3{x^2} - 4x - 7 = 0\]

OpenStudy (michele_laino):

solutions*

OpenStudy (anonymous):

okay so i need to find x

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

do i use the equation -b/2a?

OpenStudy (michele_laino):

you should get 2 solutions, namely you have to use this equation: \[\Large x = \frac{{ - b \pm \sqrt {{b^2} - 4 \times a \times c} }}{{2 \times a}}\]

rvc (rvc):

when multiplied= -21 when added = --4

OpenStudy (anonymous):

im so confused

rvc (rvc):

find two numbers such that when they are multiplied their product equals -21 and when they are added the sum equals -4

OpenStudy (anonymous):

where did you get -21?

rvc (rvc):

\[\large 3x^2-4x-7=0\] multiply the co-efficient of \(\ x^2\) and the constant

rvc (rvc):

well i do this way if you follow this you will get (x+ )(x+ )

OpenStudy (anonymous):

okay so just go step by step im confused.

OpenStudy (anonymous):

wait i got it is it (x+1)(3x-7)?

rvc (rvc):

okay so what will be the two numbers? example \[x^2+5x-6=0\] here we need two numbers such that their sum is 5 and their product is -6 so the two numbers are 6 and -1 6 X -1 = -6 6 + (-1) = 5 so the next step would be \[\large x^2-x+6x-6=0\] \[\large x(x-1)+6(x-1)=0\] \[\large (x-1)(x+6)=0\]